G3tH3Pxmpthat5Amx5t:

If r(x) = 3x – 1 and s(x) = 2x + 1, which expression is equivalent to (StartFraction r Over s EndFraction) (6)?

HelpASAP:

If r(x) = 3x – 1 and s(x) = 2x + 1, which expression is equivalent to (StartFraction r Over s EndFraction) (6)? Is this correct? lmk!

G3tH3Pxmpthat5Amx5t:

hey i think thats wrong this is the answer choices

G3tH3Pxmpthat5Amx5t:

StartFraction 3 (6) minus 1 Over 2 (6) + 1 EndFraction StartFraction (6) Over 2 (6) + 1 EndFraction StartFraction 36 minus 1 Over 26 + 1 EndFraction StartFraction (6) minus 1 Over (6) + 1 EndFraction

G3tH3Pxmpthat5Amx5t:

i cant

HelpASAP:

Hmmm I mustve miss calculated Hold up

HelpASAP:

Have you tried doing it urself buddy?

G3tH3Pxmpthat5Amx5t:

no

HelpASAP:

idk it tho i tried to work it out. I haven't leanred this stuff yet im only a girl in 10th grade lol

G3tH3Pxmpthat5Amx5t:

same

G3tH3Pxmpthat5Amx5t:

grde

HelpASAP:

(r/s)(6)= 3(6)-1/2(6)+1= 18-1/12+1=17/13 how bout that